X+yz = 40Y+xz = 11Z+xy = 10Find x, y, zPlease tell this ans sir ji/mam ji

Grade:10

1 Answers

Saurabh Koranglekar
4 years ago
Dear student

X = 10-yz

Now 40y + (10-yz)z = 10

And 11z+(10-yz) = 10

You have 2 equation 2 variables to solve, from here you can proceed

If problem still persists kindly comment we will be happy to help

Regards